Dr Westin Dentist: What is the average number of patients waiting in queue? Dr. Weston is a dentist who oversees the training of new dental students in a local clinic serving the needs of the general public. The clinic has six dental chairs, each manned by a student. Patients arrive at the rate of 20 per hour according to a Poisson distribution. The average time required for a dental check-up is 15 minutes according to a negative exponential distribution. What is the average number of patients waiting in queue?
Answer #1

This is a M/M/s system with following parameters

\lambda = 20 per hour

\mu = 60/15 = 4 per hour

s = 6   (number of servers. here server is a dental chair manned by a student)

\lambda/\mu = 20/4 = 5

Average number of patients waiting in queue, Lq = 2.938

-------------------------

Method 1: Value of Lq can be directly read from the table of multi-channel queue system , corresponding to \lambda /\mu = 5 and M=6 , as highlighted in the following table

λμ Μ LG Po λμ Μ LG % λμ M LG Po 3.4 4.3 5.2 .003 ,005 005 4.4 3.5 5.3 .019 029 .032 .033 .033 .015 026 .029 .030 .030 .030 .0

---------------------------------

Method 2: This is computed using Excel as shown below:

phpdkzqLa.png

EXCEL FORMULA:

C D E 1 per hour per hour A B A Mutliple-Server Waiting Line System 2 3 Input: 4 Arrival rate = 20 5 Service rate = 4 6 No. o